### v1

@@ Abstract
Click-fraud is commonly seen as a key vulnerability of pro-rata revenue sharing on music streaming platforms, whereas user-centric is largely immune. This paper develops a tractable non-cooperative model in which artists can purchase fraud activity that generates undetectable fake streams up to a technological limit. We show that pro-rata can be fraud-robust: when fraud technology is weak, honesty is a strict dominant strategy, and an efficient fraud-free equilibrium obtains. When fraud technology is strong, a unique fraud equilibrium arises, yet aggregate fake streams remain bounded. Although fraud is inefficient, the resulting redistribution may improve fairness in some cases. To mitigate fraud without abandoning pro-rata, we introduce a parametric weighted rule that interpolates between pro-rata and user-centric, and characterize parameter ranges that restore a fraud-free equilibrium under technology constraint. We also discuss implications of Spotify’s modernized royalty system for fraud incentives.
@@ 1 Introduction
Music streaming platforms (e.g., Spotify, Apple Music, and Deezer) charge users a subscription fee and share revenue with artists who supply music content, according to users’ streams and the platform’s revenue-sharing rule. Two rules commonly used in practice are pro-rata and user-centric. Under pro-rata, an artist receives a share equal to their streams divided by total streams; under user-centric, each user’s payment is allocated proportionally to the artists that user streams.
In the debate between these two rules, a major concern for pro-rata is click-fraud (Dimont, 2018). The reason is straightforward: if an artist can generate an unlimited number of artificial streams under pro-rata, they could capture almost all royalties. Examples include Vulfpeck’s Sleepify and Michael Smith’s bot accounts, as discussed in the literature (Dimont, 2018; Ghosh et al., 2025). Vulfpeck created an album consisting of silent tracks and asked fans to stream it while asleep. Michael Smith posted AI-generated songs and used thousands of bot accounts to boost streams. In both cases, fraudulent activity increased royalty payments.
@@
However, this does not mean that pro-rata’s vulnerability to fraud is insurmountable. As noted in Ghosh et al. (2025), these axioms impose no upper bound on streams generated by manipulated users. In practice, even without advanced fraud detection, platforms can detect and remove large volumes of suspicious streams. To capture this idea, we model a fraud technology that produces a fixed number of streams, $\lambda_{0}$, per unit of fraud activity paid by artists. The parameter $\lambda_{0}$ represents the maximum number of streams from each fake user that the platform cannot detect. Interpreted in relative terms, improvements in detection reduce the scope for subtle manipulation, while evolving fraud methods may make manipulation harder to detect.
@@
Previewing the main results, we show that honesty is a strict dominant strategy for all artists when fraud technology is not too strong, with the threshold depending on artists’ real streamshares. Even when fraud technology is strong and cannot be fully controlled, total fake streams remain bounded. Under pro-rata, an increase in total streams dilutes the value of each stream, reducing the marginal benefit from additional fake streams until it falls below the fraud cost. Moreover, fraud may lead to a fairer distribution: it benefits cheaters while reducing the income of high-streamshare honest artists by diluting stream value, potentially increasing the minimum payoff.
@@
We also discuss alternative approaches. First, we extend the analysis to a class of weighted rules that are convex combinations of pro-rata and user-centric. These rules differ in their ability to deter fraud and allow the platform to balance simplicity and robustness, maintaining efficiency under constraints on fraud technology. Second, we discuss Spotify’s recent modernization of its royalty system, focusing on a qualification policy that introduces an eligibility threshold for revenue sharing. We show that, when the qualification threshold is not too high, the policy either fails to deter fraud or harms all artists.
@@
This paper is the first to study strategic manipulation by artists on music streaming platforms using a non-cooperative game. By analyzing pro-rata’s vulnerability to fraud, the paper contributes to the ongoing debate between pro-rata and user-centric. Beyond conceptual discussions and axiomatic analyses (Dimont, 2018; Bergantiños and Moreno-Ternero, 2025; Ghosh et al., 2025), Alaei et al. (2022) and Lei (2023) are the main theoretical contributions in this area. Alaei et al. (2022) study a model with participation decisions by both users and artists. Lei (2023) shows that pro-rata can be more efficient and fairer than user-centric from the users’ perspective when music quality is endogenously chosen. In contrast to these long-run approaches that emphasize participation and quality, our paper focuses on artists’ short-run strategic manipulation, thereby enriching this literature.
@@
Contest technology is applied in analysis. Fraud activity links the pro-rata sharing rule to standard contest models with headstarts. The contest success function in our model is axiomatized by Yu (2025) through “No advantageous reallocation,” extending Amegashie (2006) from the symmetric case to a contest framework with equal starts. Franke et al. (2018) study lottery contests with linear contest technology in a setting that is more general than ours, focusing on the optimal design of headstarts to maximize effort. In our model, effort is purely socially wasteful, and the focus is on equilibrium conditions, bounds on aggregate fraud, and fairness comparison.
@@
The remainder of the paper is organized as follows. Section 2 presents the model. Section 3 analyzes the fraud-free and fraud equilibria and compares fairness. Section 4 discusses alternative solutions. Section 5 concludes.
@@ 2 Model
User-centric is immune to fraud (Bergantiños and Moreno-Ternero, 2025; Ghosh et al., 2025). Under user-centric, the net gain of artist $i$ is
@@
where a unit of fraud activity increases the cheating artist’s revenue by $\beta$ while costing $\delta$. Thus, fraud is not profitable under user-centric.
To simplify the analysis, let $\xi=\frac{\delta-\beta}{\beta}$ and $V=\frac{\lambda_{0}-\overline{\lambda}}{\xi\overline{\lambda}}$. Define $t_{i}=\frac{\lambda_{0}x_{i}}{m\overline{\lambda}}$ for all $i\in N$, so that, in the simplified model, artists choose the ratio of fake streams to total real streams rather than the level of fraud activity. Define the utility of artist $i\in N$ as $u_{i}(t)=\frac{\lambda_{0}G_{i}^{pr}(x)}{m\overline{\lambda}(\delta-\beta)}$, which yields
$u_{i}(t)=\frac{d_{i}+t_{i}}{1+\sum_{j\in N}t_{j}}V-t_{i}+d_{i}.$ (1)
The last term is constant and therefore does not affect incentives. All information in the model is public.
3 Analysis
3.1 Fraud-free equilibrium under weak fraud technology
In this royalty-sharing game, the most desirable outcome is an equilibrium in which all artists remain honest, thereby avoiding social waste. To address the vulnerability of the pro-rata to fraud, we first show that honesty is a strict dominant strategy when fraud technology is weak.
@@ 3.1 Fraud-free equilibrium under weak fraud technology
$t_{i}=0$ is a strict dominant strategy of artist $i$ if and only if $\lambda_{0}\leq\overline{\lambda}[1+\xi h(d_{i})]$, where
$h(d_{i})=\left\{\begin{array}[]{ll}\frac{1}{1-d_{i}}&\text{if $d_{i}\in[0,\frac{1}{2}];$}\\ 4d_{i}&\text{if $d_{i}\in(\frac{1}{2},1]$.}\end{array}\right$.
@@
From this expression, $\xi h(\cdot)$ represents a threshold for the deviation of fraud technology from average streaming. Intuitively, the left-hand side measures how closely a fake user resembles a real user: the smaller the deviation, the more subtle the manipulation. If the platform can detect and remove all fake users whose streaming is no greater than that of real users (i.e., LHS $\leq 0$), then fraud is not profitable and all artists remain honest.
@@
where $\delta-1$ is the price of the fraud service and $1-\beta$ is the platform’s retained share of the subscription fee. The fraud premium therefore measures the extra cost per fake user relative to the maximum surplus $\beta$ obtainable from fraud through the revenue pool.
As an immediate implication of Lemma 1, the following theorem characterizes the fraud-free equilibrium.
@@
$t^{*}=\mathbf{0}$ is the strict dominant strategy of all artists if and only if the fraud technology satisfies
$\lambda_{0}\leq\overline{\lambda}\bigg(1+\frac{\xi}{1-d_{\min}}\bigg),$
@@
Theorem 1 shows that fraud can be avoided under sufficiently weak fraud technology. We call the fraud technology weak if $\lambda_{0}\leq\overline{\lambda}(1+\frac{\xi}{1-d_{\min}})$, and strong if $\lambda_{0}>\overline{\lambda}(1+\frac{\xi}{1-d_{\min}})$.
Theorem 1 also suggests two levers for mitigating fraud. First, the platform can adjust the revenue share parameter $\beta$; in the long run, however, such changes may affect participation (Alaei et al., 2022). Second, the platform can strengthen quality checks, e.g., identity verification, to increase the cost of fraud services $\delta$ and hence the premium $\xi$. The threshold is increasing in the pivotal artist’s streamshare $d_{\min}$. However, balancing streamshares among artists can raise the threshold only up to $\overline{\lambda}(1+\frac{n}{n-1}\xi)$, which may be limited in large platforms ($n\to+\infty$).
3.2 Fraud equilibrium under strong fraud technology
@@
The upper bound of the total-to-real streams ratio $1+T^{*}$ is $V=\frac{\lambda_{0}-\overline{\lambda}}{\xi\overline{\lambda}}$.
Recall that $\frac{\lambda_{0}-\overline{\lambda}}{\overline{\lambda}}$ measures the relative advantage obtained from fraud, and $\xi$ is the fraud premium, measuring the percentage extra cost. Thus, $V$ summarizes the relative advantage from socially wasteful manipulation.
@@
3.3 Fairness comparison
So far we have studied equilibrium outcomes under weak and strong fraud technology. The fraud-free equilibrium is efficient but requires sufficiently weak fraud technology, while the fraud equilibrium features bounded social waste. The efficiency comparison is therefore clear: strong fraud technology leads to an inefficient outcome. However, efficiency is not the only criterion.
From the perspective of fairness, we show that fraud may sometimes improve the distribution of payoffs. The fairness criterion considered here is egalitarianism: social welfare depends on the agent with the lowest utility.
@@ 3.3 Fairness comparison
In the fraud equilibrium, the utilities of honest artists decrease because their revenues depend only on real streamshare, while total streams increase due to fraud. Fraud redistributes payoffs from high-share artists to low-share artists by diluting stream value. This redistribution can raise the minimum utility and thereby improve egalitarian fairness. Combining Proposition 2 with Proposition 1 yields an equivalent condition for when the fraud equilibrium is fairer.
@@ 4.1 Parametric weighted rule
$u_{i}^{pw}(t;\alpha)=\frac{d_{i}+t_{i}}{1+\sum_{j\in N}t_{j}}V\alpha-t_{i}+d_{i}\alpha+\frac{\lambda_{0}}{n\overline{\lambda}}\sum_{k\in M}\pi_{ik}(1-\alpha),$ (4)
where the last two terms do not depend on artist $i$’s choice of $t_{i}$.
@@
Under the weighted rule with parameter $\alpha$, $t^{*}=\mathbf{0}$ is the strict dominant strategy of all artists if and only if $\lambda_{0}\leq\overline{\lambda}[1+\frac{\xi}{\alpha(1-d_{\min})}]$.
@@ 4.2 Spotify’s modernized royalty system
We argue that this policy may strictly increase fake streams and hurt all artists. To illustrate this possibility, we focus on a reasonable range of thresholds $\widehat{d}$ that are not too high: an overly demanding threshold could affect artists who rely on streaming revenue, and a full characterization for arbitrary thresholds is beyond the scope of this paper. Within the range we consider, dishonest artists in the baseline equilibrium still cheat.
@@
Suppose instead that the platform chooses a threshold $\widehat{d}$ slightly above $d^{*}$. [^2] The range of slightly-high $\widehat{d}$ is non-empty, as shown in Appendix A.8. In this range, all artists cheat at least the same as in the baseline equilibrium, and no artist is eliminated. Then a low-streamshare artist chooses either $t_{i}=0$ or $t_{i}\geq\widehat{d}-d_{i}$. In a new equilibrium where all initial cheaters still cheat, their best response is the just-qualify strategy $\widehat{d}-d_{i}$, because the resulting increase in total fake streamshare reduces the marginal return to additional fraud. Let $\widehat{t}$ be defined by $\widehat{t}_{i}=\max\{\widehat{d}-d_{i},0\}$ for all $i\in N$, and $\widehat{T}$ denote aggregate fraud under $\widehat{t}$. The strategy profile $\widehat{t}$ is the equilibrium under threshold $\widehat{d}$. In this equilibrium, artist $i$ with $\widehat{d}>d_{i}$ will cheat more, i.e., $\widehat{t}_{i}>t_{i}^{*}$; and artist $i$ with $\widehat{d}\leq d_{i}$ will keep honest, i.e., $\widehat{t}_{i}=t^{*}$, and thus $\widehat{T}-\widehat{t}_{i}>T^{*}-t_{i}^{*}$. Consequently, for all $i\in N$,
[…]

### v2

@@ Abstract
Click-fraud is commonly seen as a key vulnerability of pro-rata revenue sharing rule on music streaming platforms, whereas user-centric is largely immune. This paper develops a tractable non-cooperative model in which artists can purchase fraud activity that generates undetectable fake streams up to a technological limit. We defend pro-rata by showing that it is fraud-robust: when fraud technology is weak, honesty is a strictly dominant strategy, and an efficient fraud-free equilibrium obtains; when fraud technology is strong, a unique fraud equilibrium arises, yet aggregate fake streams remain bounded. Although fraud is inefficient, the resulting redistribution may improve fairness in some cases. To mitigate fraud without abandoning pro-rata, we introduce a parametric weighted rule that interpolates between pro-rata and user-centric, and characterize parameter ranges that restore a fraud-free equilibrium under technology constraint. We also discuss implications of Spotify’s modernized royalty system for fraud incentives.
@@ 1 Introduction
Music streaming platforms (e.g., Spotify, Apple Music, and Deezer) charge users a subscription fee and share revenue with artists who supply music content. Royalties are typically allocated as a function of users’ streams and the platform’s revenue-sharing rule. Two rules commonly used in practice are pro-rata and user-centric. Under pro-rata, an artist receives a share equal to their streams divided by total streams; under user-centric, each user’s payment is allocated proportionally to the artists that user streams.
There is an ongoing debate about which rule should be used, especially between pro-rata and user-centric. At the center of this debate, Dimont (2018) raises several major concerns about pro-rata, and advocates a pay-per-subscriber model, such as user-centric. [^1] In addition to the problem of disproportionate royalties and click-fraud mentioned in this paper, Dimont (2018) also raises a concern about the silence of major labels. We omit this here because it is unrelated to the pro-rata mechanism itself. One concern is the problem of disproportionate royalties. The argument highlights cross-subsidization from low-usage users to high-usage users, as well as the induced cross-subsidization between artists listened to by heavy users and those listened to by light users. Alaei et al. (2022) defend pro-rata on this point, showing that despite cross-subsidization across users, pro-rata can be preferred by both artists and the platform.
Another major concern is click-fraud. The intuition is straightforward: if an artist can generate an unlimited number of artificial streams under pro-rata, they could capture nearly all royalties. Real-world examples include Vulfpeck’s Sleepify and Michael Smith’s bot accounts, as discussed in the literature (Dimont, 2018; Ghosh et al., 2025). Vulfpeck created an album consisting of silent tracks and asked fans to stream it while asleep. Michael Smith, who was ultimately arrested, uploaded AI-generated songs and used thousands of bot accounts to boost streams. In both cases, fraudulent activity increased royalty payments.
@@
However, this does not mean that pro-rata’s vulnerability to fraud is insurmountable. As noted in Ghosh et al. (2025), these axioms impose no upper bound on streams generated by manipulated users. In practice, even without sophisticated fraud detection, platforms can detect and remove large volumes of suspicious streams. To capture this idea, we model a fraud technology that produces a fixed number of streams, $\lambda_{0}$, per unit of fraud activity purchased by artists. The parameter $\lambda_{0}$ represents the maximum number of streams from each fake user that the platform fails to detect. Interpreted in relative terms, improved detection reduces the scope for subtle manipulation, while evolving fraud methods may make manipulation harder to detect.
@@
This paper defends pro-rata with respect to click-fraud. Previewing the main results, we show that pro-rata is fraud-robust: honesty is a strictly dominant strategy for all artists when fraud technology is not too strong, with a threshold that depends on artists’ real streamshares. Even when fraud technology is strong and cannot be fully controlled, total fake streams remain bounded. Under pro-rata, an increase in total streams dilutes the value of each stream, reducing the marginal benefit of additional fake streams until it falls below the fraud cost. Moreover, fraud may lead to a fairer distribution: it benefits cheaters while reducing the income of high-streamshare honest artists by diluting stream value, potentially increasing the minimum payoff. Thus, the problem of disproportionate royalties may also be mitigated under pro-rata due to click-fraud.
@@
We also discuss alternative approaches. First, we extend the analysis to a class of weighted rules that are convex combinations of pro-rata and user-centric. These rules differ in their ability to deter fraud and allow the platform to balance simplicity and robustness while maintaining efficiency under constraints on fraud technology. Second, we discuss Spotify’s recent modernization of its royalty system, focusing on a qualification policy that introduces an eligibility threshold for revenue sharing. We show that when the qualification threshold is not too high, the policy either fails to deter fraud or harms all artists.
@@
This paper is the first to study strategic manipulation by artists on music streaming platforms using a non-cooperative game. By analyzing pro-rata’s vulnerability to fraud, the paper contributes to the ongoing debate between pro-rata and user-centric. Beyond conceptual discussions and axiomatic analyses (Dimont, 2018; Bergantiños and Moreno-Ternero, 2025; Ghosh et al., 2025), Alaei et al. (2022) and Lei (2023) are the main theoretical contributions in this area. Alaei et al. (2022) study a model with participation decisions by both users and artists. Lei (2023) shows that pro-rata can be more efficient and fairer than user-centric from users’ perspectives when music quality is endogenously chosen. In contrast to these long-run approaches that emphasize participation and quality, our paper focuses on artists’ short-run strategic manipulation, thereby enriching this literature.
@@
Contest technology is applied in the analysis. Fraud activity links the pro-rata revenue-sharing rule to standard contest models with head starts. The contest success function in our model is axiomatized by Yu (2026) through “No advantageous reallocation,” extending Amegashie (2006) from the symmetric case to a contest framework with equal starts. Franke et al. (2018) study lottery contests with linear contest technology in a setting more general than ours, focusing on the optimal design of headstarts to maximize effort. In our model, efforts are purely socially wasteful, and the focus is on equilibrium conditions, bounds on aggregate fraud, and fairness comparisons.
@@
The remainder of the paper is organized as follows. Section 2 presents the model. Section 3 characterizes the fraud-free equilibrium under weak fraud technology. Section 4 analyzes the fraud equilibrium under strong fraud technology. Section 5 compares the fairness driven by fraud. Section 6 discusses alternative solutions. Section 7 concludes.
@@ 2 Model
Remark that user-centric is immune to fraud (Bergantiños and Moreno-Ternero, 2025; Ghosh et al., 2025). Under user-centric, the net gain of artist $i$ is
@@
where a unit of fraud activity increases the cheating artist’s revenue by $\beta$ while costing $\delta$. That is, fraud is not profitable under user-centric.
To simplify the analysis, let $\xi=\frac{\delta-\beta}{\beta}$ and $V=\frac{\lambda_{0}-\overline{\lambda}}{\xi\overline{\lambda}}$. Define the ratio of fake streams on artist $i$ to total real streams $t_{i}=\frac{\lambda_{0}x_{i}}{m\overline{\lambda}}$ for all $i\in N$. In the simplified model, each artist $i$ choose strategy $t_{i}\in\mathbb{R}_{+}$. Define the utility of artist $i\in N$ as $u_{i}(t)=\frac{\lambda_{0}G_{i}^{pr}(x)}{m\overline{\lambda}(\delta-\beta)}$, which yields
$u_{i}(t)=\frac{d_{i}+t_{i}}{1+\sum_{j\in N}t_{j}}V-t_{i}+\frac{d_{i}}{\xi}.$ (1)
The last term is constant and therefore does not affect incentives. All information in the model is public. In the following analysis, we focus on dominant strategy equilibrium when the fraud technology is weak, and on Nash equilibrium otherwise.
3 Case 1: Weak fraud technology
In this paper, efficiency is measured by the total surplus of artists under a utilitarian criterion. We call an equilibrium efficient if it maximizes the summation of artists’ utilities. Note that each unit of fraud activity adds a taxed subscription fee $\beta$ to the revenue pool shared by artists, but incurs a fraud cost $\delta$ for the cheating artist. Thus, fraud is socially wasteful: it reduces total surplus by $\delta-\beta>0$ per unit of fraud activity. Consequently, the most desirable outcome would be an equilibrium in which all artists remain honest, thereby avoiding social waste.
To illustrate the possibility of this desirable outcome, we first show that honesty is a strictly dominant strategy when fraud technology is weak.
@@ 3 Case 1: Weak fraud technology
$t_{i}=0$ is a strictly dominant strategy of artist $i$ if and only if $\lambda_{0}\leq\overline{\lambda}[1+\xi h(d_{i})]$, where
$h(d_{i})=\left\{\begin{array}[]{ll}\frac{1}{1-d_{i}}&\mathrm{if}\ d_{i}\in[0,\frac{1}{2}];\\ 4d_{i}&\mathrm{if}\ d_{i}\in(\frac{1}{2},1].\end{array}\right$.
@@
From this expression, $\xi h(\cdot)$ represents a threshold for the deviation of fraud technology from average streaming. Intuitively, the left-hand side (LHS) measures how closely a fake user resembles a real user: the smaller the deviation, the more subtle the manipulation. If the platform can detect and remove all fake users whose streaming is no greater than that of real users (i.e., LHS $\leq 0$), then fraud is not profitable and all artists remain honest.
@@
where $\delta-1$ is the price of the fraud service and $1-\beta$ is the platform’s retained share of the subscription fee. The fraud premium measures the extra cost per fake user relative to the maximum surplus $\beta$ obtainable from fraud through the revenue pool.
We refer to an equilibrium $t^{*}$ as a fraud-free equilibrium if there is no fake stream generated, i.e., $t^{*}=\mathbf{0}$. As an immediate implication of Lemma 1, the following theorem characterizes the fraud-free equilibrium.
@@
$t^{*}=\mathbf{0}$ is the dominant strategy equilibrium if and only if the fraud technology satisfies
$\lambda_{0}\leq\overline{\lambda}\bigg(1+\frac{\xi}{1-d_{\min}}\bigg),$ (2)
@@
We call the fraud technology weak if (2) holds, and strong otherwise. Theorem 1 shows that fraud can be deterred under a sufficiently weak fraud technology. Hence a unique fraud-free equilibrium exists. This fraud-free equilibrium is efficient because it involves no waste. In this equilibrium, honesty is a strictly dominant strategy for each artist, aligning the incentive with the utilitarian objective.
Theorem 1 also suggests two levers for mitigating fraud. First, the platform can adjust the revenue share parameter $\beta$; in the long run, however, such changes may affect participation (Alaei et al., 2022). Second, the platform can strengthen quality checks, e.g., identity verification, to increase the cost of fraud services $\delta$ and hence the premium $\xi$. The threshold is increasing in the pivotal artist’s streamshare $d_{\min}$. However, balancing streamshares among artists can raise the threshold only up to $\overline{\lambda}(1+\frac{n}{n-1}\xi)$, which may be limited in a large platform ($n\to+\infty$).
4 Case 2: Strong fraud technology
@@
The upper bound of the total-to-real streams ratio $1+T^{*}$ is $V$.
Recall that $\frac{\lambda_{0}-\overline{\lambda}}{\overline{\lambda}}$ measures the relative advantage obtained from fraud, and $\xi$ is the fraud premium, measuring the percentage extra cost. Thus, $V=\frac{\lambda_{0}-\overline{\lambda}}{\xi\overline{\lambda}}$ summarizes the relative advantage from socially wasteful manipulation.
@@
5 Fairness comparison
So far we have studied equilibrium outcomes under weak and strong fraud technologies. The fraud-free equilibrium is efficient but requires sufficiently weak fraud technology, while the fraud equilibrium features bounded social waste. The efficiency comparison is therefore clear: strong fraud technology leads to an inefficient outcome.
However, efficiency is not the only criterion. From the perspective of fairness, we show that fraud may sometimes improve the distribution of payoffs. The fairness criterion considered here is egalitarianism: social welfare depends on the agent with the lowest utility.
@@ 5 Fairness comparison
In the fraud equilibrium, the utilities of honest artists decrease because their revenues depend only on real streamshare, while total streams increase due to fraud. Fraud redistributes payoffs from high-streamshare artists to low-streamshare artists by diluting stream value. This redistribution can raise the minimum utility and thereby improve egalitarian fairness. Combining Proposition 2 with Proposition 1 yields an equivalent condition for when the fraud equilibrium is fairer.
@@ 6.1 Parametric weighted rule
$u_{i}^{pw}(t;\alpha)=\frac{d_{i}+t_{i}}{1+\sum_{j\in N}t_{j}}V\alpha-t_{i}+\frac{1}{\xi}\Bigg[d_{i}\alpha+\frac{\lambda_{0}}{m\overline{\lambda}}\sum_{k\in M}\pi_{ik}(1-\alpha)\Bigg],$ (5)
where the last term does not depend on artist $i$’s choice of $t_{i}$.
@@
Under the weighted rule with parameter $\alpha$, $t^{*}=\mathbf{0}$ is the strictly dominant strategy of all artists if and only if $\lambda_{0}\leq\overline{\lambda}[1+\frac{\xi}{\alpha(1-d_{\min})}]$.
@@ 6.2 Spotify’s modernized royalty system
We argue that this policy may strictly increase fake streams and harm all artists. To illustrate this possibility, we focus on a reasonable range of thresholds $\widehat{d}$ that are not too high: an overly demanding threshold could affect artists who rely on streaming revenue, and a full characterization for arbitrary thresholds is beyond the scope of this paper. Within the range we consider, dishonest artists in the baseline equilibrium still cheat.
@@
Suppose instead that the platform chooses a threshold $\widehat{d}$ slightly above $d^{*}$. [^3] The range of slightly high $\widehat{d}$ is non-empty, as shown in Appendix A.8. In this range, all artists cheat at least the same as in the baseline equilibrium, and no artist is eliminated. Then a low-streamshare artist chooses either $t_{i}=0$ or $t_{i}\geq\widehat{d}-d_{i}$. In a new equilibrium where all initial cheaters still cheat, their best response is the just-qualify strategy $\widehat{d}-d_{i}$, because the resulting increase in total fake streamshare reduces the marginal return to additional fraud. Let $\widehat{t}$ be defined by $\widehat{t}_{i}=\max\{\widehat{d}-d_{i},0\}$ for all $i\in N$, and $\widehat{T}$ denote aggregate fraud under $\widehat{t}$. The strategy profile $\widehat{t}$ is the equilibrium under threshold $\widehat{d}$. In this equilibrium, artist $i$ with $\widehat{d}>d_{i}$ will cheat more, i.e., $\widehat{t}_{i}>t_{i}^{*}$; and artist $i$ with $\widehat{d}\leq d_{i}$ will keep honest, i.e., $\widehat{t}_{i}=t^{*}$, and thus $\widehat{T}-\widehat{t}_{i}>T^{*}-t_{i}^{*}$. Consequently, for all $i\in N$,
$u_{i}(\widehat{t})=\frac{d_{i}+\widehat{t}_{i}}{1+\widehat{T}}V-\widehat{t}_{i}+\frac{d_{i}}{\xi}\leq\frac{d_{i}+\widehat{t}_{i}}{1+\widehat{t}_{i}+T^{*}-t_{i}^{*}}V-\widehat{t}_{i}+\frac{d_{i}}{\xi}=u_{i}(\widehat{t}_{i},t_{-i}^{*})\leq u_{i}(t^{*}),$